OpenStudy (anonymous):

Two spheres of the same density have a ratio of 4 to 9 in surface area. If the small sphere weighs 10kg, what does the sphere weigh

ganeshie8 (ganeshie8):

surface are of sphere is proportional to r^2

ganeshie8 (ganeshie8):

\(\huge r1^2 : r2^2 = 4 : 9\)

OpenStudy (anonymous):

it says the answer is 33.75 kg but i don't know how to get that

ganeshie8 (ganeshie8):

yep.. il walk you through...

OpenStudy (anonymous):

okay. Thanks

ganeshie8 (ganeshie8):

\(\huge r1 : r2 = 2 : 3 \)

ganeshie8 (ganeshie8):

volume of sphere is proportional to r^3 right ?

OpenStudy (anonymous):

i guess

ganeshie8 (ganeshie8):

yup it is.. we dont need the actual formula here, as we are dealing with ratios. ..

ganeshie8 (ganeshie8):

\(\huge r1^3 : r2^3 = 8 : 27\)

ganeshie8 (ganeshie8):

If the small sphere weighs 10kg, what does the sphere weigh

OpenStudy (anonymous):

what does r1^3 and r2^3 mean?

ganeshie8 (ganeshie8):

sphere with \(r1\) radius weighs \(10kg\) we need to find the weight of sphere with \(r2\) radius

ganeshie8 (ganeshie8):

they are volume ratios. since the density is same for both spheres, we can use the volume ratio.

ganeshie8 (ganeshie8):

volume of sphere = \(\frac{4}{3}\pi r^3\)

ganeshie8 (ganeshie8):

is it clear ?

OpenStudy (anonymous):

i don't really get that

ganeshie8 (ganeshie8):

in question, we are given small sphere's weight. and asking us to compute large sphere's weight right ?

OpenStudy (anonymous):

yes

ganeshie8 (ganeshie8):

1) weight of a sphere is proportional to volume 2) volume of sphere = 4/3 pi r^3 3) so we can take ratio of r^3 to the weight

ganeshie8 (ganeshie8):

since they are of same density, weight is proportional to volume. so we can compare volume ratio to weights, right ?

OpenStudy (anonymous):

yea

ganeshie8 (ganeshie8):

and volume is proportional to r^3

ganeshie8 (ganeshie8):

thats how we got to 8:27

ganeshie8 (ganeshie8):

we are given surface area ration : 4 : 9

ganeshie8 (ganeshie8):

we started from above, right ?

ganeshie8 (ganeshie8):

surface area is proportional to r^2

ganeshie8 (ganeshie8):

so, ratio of radius must be : 2 : 3 so, ration of volume must be 2^3 : 3^3

OpenStudy (anonymous):

okay

ganeshie8 (ganeshie8):

8:27 = 10 : x x = 27*10/8

ganeshie8 (ganeshie8):

we are equating volume ratios to weight ratios, because density is same

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

oh I got the answer

ganeshie8 (ganeshie8):

below info is needed for this problem : surface area of sphere = \(4\pi r ^2\) volume of sphere = \(\frac{4}{3}\pi r^3\)

ganeshie8 (ganeshie8):

since weight follows volume, we took the volume ratio.

ganeshie8 (ganeshie8):

hope it becomes clear later :)... maybe !!

OpenStudy (anonymous):

thank you for your help

ganeshie8 (ganeshie8):

u welcome :)